# CDKFormer: Contextual Deviation Knowledge-Based Transformer for Long-Tail Trajectory Prediction

### v1

@@ Abstract
Predicting the future movements of surrounding vehicles is essential for ensuring the safe operation and efficient navigation of autonomous vehicles (AVs) in urban traffic environments. Existing vehicle trajectory prediction methods primarily focus on improving overall performance, yet they struggle to address long-tail scenarios effectively. This limitation often leads to poor predictions in rare cases, significantly increasing the risk of safety incidents. Taking Argoverse 2 motion forecasting dataset as an example, we first investigate the long-tail characteristics in trajectory samples from two perspectives, individual motion and group interaction, and deriving deviation features to distinguish abnormal from regular scenarios. On this basis, we propose CDKFormer, a Contextual Deviation Knowledge-based Transformer model for long-tail trajectory prediction. CDKFormer integrates an attention-based scene context fusion module to encode spatiotemporal interaction and road topology. An additional deviation feature fusion module is proposed to capture the dynamic deviations in the target vehicle’s status. We further introduce a dual query-based decoder, supported by a multistream decoder block, to sequentially decode heterogeneous scene deviation features and generate multimodal trajectory predictions. Extensive experiments demonstrate that CDKFormer achieves state-of-the-art performance, significantly enhancing prediction accuracy and robustness for long-tailed trajectories compared to existing methods, thus advancing the reliability of AVs in complex real-world environments.
Index Terms:
@@
trajectory prediction; long-tail learning; Transformer; contextual deviation knowledge; query-based decoding
I Introduction
@@
Autonomous vehicles rely on trajectory prediction to anticipate the future movements of other traffic participants [1]. This predictive ability is crucial for enabling well-informed driving decisions that prioritize both safety and traffic efficiency [2, 3, 4]. Current state-of-the-art (SOTA) trajectory prediction methods have demonstrated impressive performance on large-scale datasets such as Waymo Open Motion Dataset (WOMD) [5], Argoverse [6] and nuScenes [7] motion forecasting datasets. However, real-world traffic scenarios exhibit a long-tailed distribution, characterized by a multitude of rare and unusual events—such as sudden stops, erratic lane changes, or unexpected obstacles—that occur infrequently, while common and predictable scenarios, like straight driving and lane-following, dominate. This phenomenon is often referred to as the ”curse of rarity” [8]. Such an imbalance causes deep learning-based vehicle trajectory prediction models to become biased toward these frequent scenarios, struggling to accurately predict rare and atypical events that are underrepresented in the training data. Consequently, this undermines the models’ robustness and reliability, posing significant challenges for trajectory prediction models in safety-critical situations [9].
@@
Addressing the learning of trajectory prediction models on imbalanced datasets is therefore of vital importance to enhance the reliability and safety of autonomous driving systems. Current approaches to long-tail trajectory prediction employ techniques such as data augmentation [10, 11], loss design [12], contrastive learning [13, 14], and mixture of experts [15] to improve performance on tail samples. Existing methods, however, typically overlook the explicit consideration of factors that cause long-tailed samples to be both rare and challenging to predict. A significant challenge lies in the underrepresentation of features that characterize long-tail traffic scenarios. Consequently, it becomes imperative to identify measures capable of effectively characterizing and distinguishing long-tail scenarios from common ones, and accordingly design proper methods to fuse these measures in the model.
@@
In this study, we first conduct a comprehensive analysis of the characteristics of tailed samples and derive key metrics that quantify the target vehicle’s deviations from typical motion patterns and interactions with other agents, collectively termed deviation features. On this basis, we propose CDKFormer, a Contextual Deviation Knowledge-based Transformer model for long-tail trajectory prediction. CDKFormer is designed to learn deviation features and scene contextual features using Transformer architectures. These features are subsequently decoded using mode queries and dual future queries to generate robust trajectory predictions. We believe this will allow our model to actively focus on the abnormal parts of the traffic environment and learn more robust representations of both common and uncommon traffic conditions.
@@
  • We develop a dual query-based decoder to generate multimodal trajectory predictions. Supported by a multistream decoder block, the decoder decodes heterogeneous scene deviation features sequentially.
@@
The remainder of the paper is organized as follows. Section II reviews the related work. Section III defines the problem and explores the characteristics of long-tailed trajectory samples. Section IV describes the proposed model in detail. Section V presents the experimental results. Section VI concludes this work and outlines possible future directions.
II Literature Review
II-A Vehicle Trajectory Prediction
@@
Vehicle trajectory prediction aims to forecast the future movements of vehicles in dynamic traffic environments. Extensive research has been dedicated to this field, leading to the development of various models that consider both vehicle dynamics and interactions among agents and map elements.
@@
A major challenge in vehicle trajectory prediction lies in understanding agent interaction patterns. Early approaches used physics-based models [16, 17] or maneuver-based motion models [18] to estimate future trajectories. Recent advances [19, 20, 21, 22, 23, 24] have leveraged deep learning models to address this challenge, enabling models to better capture the interaction patterns between vehicles and their surroundings, such as the road network and nearby objects, to improve prediction accuracy. Some approaches [25, 26, 27] further emphasize the integration of map features as vectors to enhance trajectory prediction performance. Encoding these map features in the form of polylines has proven particularly effective.
@@
Transformers have also shown great promise in trajectory prediction tasks due to their effectiveness in processing long-range sequential data and modeling complex interactions through attention mechanisms [28, 20, 21, 29]. For instance, QCNet [21] leverages attention mechanisms to model interactions at different spatial and temporal scales, enabling more accurate and robust predictions in real-time traffic scenarios. Furthermore, some studies explore variations in Transformer architectures [30] or attention mechanism [31] to better fuse spatio-temporal contextual information.
II-B Query-Based Trajectory Decoding
@@
Most recent end-to-end trajectory prediction models adopt a query-based trajectory decoding paradigm, inspired by Detection Transformer (DETR) [32] from the object detection field. Query refers to a set of learnable embedding vectors that serve as placeholders for future trajectory predictions. These queries interact with encoded scene features through attention mechanisms within the Transformer architecture, enabling the model to generate distinct and contextually relevant trajectory forecasts. Various terms have been used to describe this underlying concept, such as ”proposals” in mmTransformer [28], ”anchor trajectories” in MultiPath++, ”queries” in MTR [20], SEPT [33], and QCNet [21], etc.
@@
Various studies have contributed to improvements in query design. Early studies ultilize predefined queries to inform the model of possible endpoints [34, 35, 20]. Recent studies purpose to decode trajctories dynamically with learnable queries [21, 36]. The queries are designed to capture the learned contextual information with cross-attention mechanism and produce multi-modal future trajectories. A recent study [2] proposes endpoint-risk-combined intention queries as prediction priors to support risk-aware risk prediction.
II-C Long-Tail Trajectory Prediction
@@
Long-tail learning addresses the challenge of imbalanced data distributions, where a large portion of the dataset consists of rare or less frequent examples. In the trajectory prediction field, various strategies have been proposed to tackle this issue, including data augmentation[10, 11], loss design [12], contrastive learning [13, 14], mixture of experts [15], model ensemble [37], etc.
@@
Input data augmentation techniques are utilized to improve the accuracy and robustness of trajectory prediction, incorporating strategies such as heading rotation, scene flipping, and adding random noise. These strategies have been shown to increase robustness against adversarial patterns in trajectories [38, 10]. The design of synthetic driving data has also demonstrated notable benefits for trajectory prediction [11].
@@
Contrastive learning (CL) was first proposed to deal with long-tail issuse in trajectory prediction by [13]. They propose to improve long-tail trajectory prediction performance by forcing similar samples together and pulling dissimilar samples apart in the feature space by a contrastive loss function. [14] propose FEND, a feature-enhanced distribution aware CL framework that ultlize prototypical contrastive learning (PCL). A following study [39] further integrates contextual scene information in the contrastive learning framework. However, the scene contextual interaction information is not explicitly considered in their contrasitive learning framework. Researchers [40] also extends the use of contrastive learning by considering subclasses dynamically. [41] propose a hierarchical wave-semantic contrastive learning (Hi-SCL) framework, which maintains a collection of feature-enhanced hierarchical prototypes, dynamically steering trajectory samples closer or pushing them farther away.
@@
Contrastive learning typically functions by differentiating between positive and negative sample pairs. In current contrastive learning-based trajectory prediction methods, the contrastive sample pairs are often constructed based on either the difficulty scores [13] or the clustering of focal agent motion patterns [14], frequently overlooking information from interacting agents and the dynamic environment. This approach poses significant challenge in effectively constructing positive and negative sample pairs that incorporate scene semantics [42], which in turn hinders the development of a robust contrastive loss function. In this paper, we avoid the challenge of attempting to apply contrastive learning to consistently consider both interacting- and scene-level clustering. Alternatively, we approach the long-tail trajectory prediction through deviation knowledge fusion and dual query-based decoding with loss reweighting. We posit that both the individual motion and interaction patterns contribute to the rarity of traffic scenarios and should be carefully designed and incorporated into the model design.
III Long-Tail Characteristics of Trajectory Data
III-A Preliminaries
@@
Given a series of $d_{a}$-dimensional observed motion features $\bm{X}\in\mathbb{R}^{N_{a}\times T_{o}\times d_{a}}$ for $N_{a}$ agents, including the target vehicle and its surrounding agents, over a time span $T_{o}$ and high-definition (HD) map vectors $\bm{I}\in\mathbb{R}^{N_{m}\times l_{m}\times d_{m}}$ of $N_{m}$ map polylines, our objective is to predict future positions $\bm{Y}\in\mathbb{R}^{T_{f}\times 2}$ of the target vehicle in a certain time horizon $T_{f}$. Specifically, we seek to train a neural network to model the mapping $f(\bm{Y}|\bm{X},\bm{I})$. Additionally, we introduce a long-tail score $S$ to quantify the likelihood of a sample belonging to the tail of the data distribution. On this basis, we seek to enhance the model’s predictive performance in these long-tail scenarios while not affecting the overall performance.
III-B Long-Tail Definition
@@
Long-tail learning aims to train a deep neural network on a dataset with a long-tailed class distribution, where a small proportion of classes contain a large number of samples, while the majority of classes are represented by only a few samples [43]. In classification tasks, it is relatively straightforward to identify tail samples. However, long-tail trajectory prediction is essentially a long-tail regression task [44], where the labels (future trajectories) are continuous values. In this case, hard classification boundaries among classes do not exist [43]. Previous studies identify tailed samples by a difficulty score, which is computed by the (final) displacement error performance of a Kalman filter [45, 13] or a trained prediction network [14].
@@
In this study, we propose measuring a tailed sample with both its difficulty and rarity. We believe that tailed trajectories are not only samples that are challenging to predict, but also tend to exhibit more complex motion patterns and diverse interaction types, making them relatively rare in spatio-temporal distribution. While these two aspects are not identical, they are closely related and should be considered together in a holistic manner. Similar to [14], we quantify the difficulty score $S_{d}$ as the average displacement error of the predictions of a trained model. The rarity score $S_{r}$ is determined by the negative log-likelihood of a Gaussian Mixture Model (GMM). As illustrated in Fig. 2, the GMM models the distribution of trajectory final points. Samples that are less likely to occur within the learned distribution (i.e., those farther from the Gaussian components) are assigned higher rarity scores. Finally, the tail score $S$ is computed as the square root product of difficulty and rarity:
@@
The tail score distribution of training samples in Argoverse 2 motion forecasting dataset is demonstrated in Fig. 2.
Figure 1: Rarity score distribution of the training samples in Argoverse 2 motion forecasting dataset. The normalized trajectory endpoints are fitted to a GMM with 6 components. For each trajectory sample, the rarity score is calculated as the negative log-likelihood of its endpoint under the learned distribution. These scores are then normalized to the range [0, 1], with higher scores indicating higher rarity.
@@
III-C Long-Tail Characteristics

### v2

@@ Abstract
Predicting the future movements of surrounding vehicles is essential for ensuring the safe operation and efficient navigation of autonomous vehicles (AVs) in urban traffic environments. Existing vehicle trajectory prediction methods focus primarily on improving overall performance, yet they struggle to address long-tail scenarios effectively. This limitation often leads to poor predictions in rare cases, significantly increasing the risk of safety incidents. Taking Argoverse 2 motion forecasting dataset as an example, we first investigate the long-tail characteristics in trajectory samples from two perspectives, individual motion and group interaction, and deriving deviation features to distinguish abnormal from regular scenarios. On this basis, we propose CDKFormer, a contextual deviation knowledge-based Transformer model for long-tail trajectory prediction. CDKFormer integrates an attention-based scene context fusion module to encode spatiotemporal interaction and road topology. An additional deviation feature fusion module is proposed to capture dynamic deviations in the target vehicle’s status. We further introduce a dual query-based decoder, supported by a multistream decoder block, to sequentially decode heterogeneous scene deviation features and generate multimodal trajectory predictions. Extensive experiments demonstrate that CDKFormer achieves state-of-the-art performance, significantly enhancing prediction accuracy and robustness for long-tailed trajectories compared to existing methods, thus advancing the reliability of AVs in complex real-world environments.
@@
Keywords: trajectory prediction; long-tail learning; Transformer; contextual deviation knowledge; query-based decoding
1 Introduction
@@
Autonomous vehicles rely on trajectory prediction to anticipate the future movements of other traffic participants [huang2022survey]. This predictive ability is crucial for enabling well-informed driving decisions that prioritize both safety and traffic efficiency [zhao2025multi, yang2024towards, liu2024integrated]. Current state-of-the-art (SOTA) trajectory prediction methods have demonstrated impressive performance on large-scale datasets such as Waymo Open Motion Dataset (WOMD) [ettinger2021large], Argoverse [wilson2023argoverse] and nuScenes [caesar2020nuscenes] motion forecasting datasets. However, real-world traffic scenarios exhibit a long-tailed distribution, characterized by a multitude of rare and unusual events—such as sudden stops, erratic lane changes, or unexpected obstacles—that occur infrequently, while common and predictable scenarios, like straight driving and lane-following, dominate. This phenomenon is often referred to as the ”curse of rarity” [liu2024curse]. Such an imbalance causes deep learning-based vehicle trajectory prediction models to become biased toward these frequent scenarios, struggling to accurately predict rare and atypical events that are underrepresented in the training data. Consequently, this undermines the models’ robustness and reliability, posing significant challenges for trajectory prediction models in safety-critical situations [ding2023survey].
@@
Addressing the learning of trajectory prediction models on imbalanced datasets is therefore of vital importance to enhance the reliability and safety of autonomous driving systems. Current approaches to long-tail trajectory prediction employ techniques such as data augmentation [bahari2022vehicle, li2024pre], loss design [kozerawski2022taming], contrastive learning [makansi2021exposing, wang2023fend], and mixture of experts [mercurius2024amend] to improve performance on tail samples. Existing methods, however, typically overlook the explicit consideration of factors that cause long-tail samples to be both rare and challenging to predict. A significant challenge lies in the underrepresentation of features that characterize long-tail traffic scenarios. Consequently, it becomes imperative to identify measures capable of effectively characterizing and distinguishing long-tail scenarios from common ones, and accordingly design proper methods to fuse these measures in the model.
@@
In this study, we first conduct a comprehensive analysis of the characteristics of tail samples and derive key metrics that quantify the target vehicle’s deviations from typical motion patterns and interactions with other agents, collectively termed deviation features. On this basis, we propose CDKFormer, a contextual deviation knowledge-based Transformer model for long-tail trajectory prediction. CDKFormer is designed to learn deviation features and scene contextual features using Transformer architectures. These features are subsequently decoded using mode queries and dual future queries to generate robust trajectory predictions. We believe this will allow our model to actively focus on the abnormal parts of the traffic environment and learn more robust representations of both common and uncommon traffic conditions.
@@
  • We develop a dual query-based decoder to generate multimodal trajectory predictions. Supported by a multistream decoder block, the decoder sequentially decodes heterogeneous scene deviation features.
@@
The remainder of the paper is organized as follows. Section 2 reviews the related work. Section 3 defines the problem and explores the characteristics of long-tailed trajectory samples. Section 4 describes the proposed model in detail. Section 5 presents the experimental results. Section 6 concludes this work and outlines possible future directions.
2 Literature Review
2.1 Vehicle Trajectory Prediction
@@
Vehicle trajectory prediction aims to forecast future movements of vehicles in dynamic traffic environments. Extensive research has been dedicated to this field, leading to the development of various models that consider both vehicle dynamics and interactions among agents and map elements.
@@
A major challenge in vehicle trajectory prediction lies in understanding agent interaction patterns. The early approaches used physics-based models [lin2000vehicle, polychronopoulos2007sensor] or maneuver-based motion models [gindele2010probabilistic] to estimate future trajectories. Recent advances [salzmann2020trajectron++, shi2022motion, zhou2023query, geng2023physics, zhao2025multi, yang2025interactive] have leveraged deep learning models to address this challenge, allowing models to better capture the interaction patterns between vehicles and their surroundings, such as the road network and nearby objects, to improve prediction accuracy. Some approaches [liang2020learning, gao2020vectornet, liu2024laformer] further emphasize the integration of map features as vectors to enhance trajectory prediction performance. Encoding these map features in the form of polylines has proven particularly effective.
@@
Transformers have also shown great promise in trajectory prediction tasks due to their effectiveness in processing long-range sequential data and modeling complex interactions through attention mechanisms [liu2021multimodal, shi2022motion, zhou2023query, yu2020spatio, zhang2022explainable, geng2023dynamic]. For instance, QCNet [zhou2023query] leverages attention mechanisms to model interactions on different spatial and temporal scales, enabling more accurate and robust predictions in real-time traffic scenarios. Furthermore, some studies explore variations in Transformer architectures [lian2024hierarchical] or attention mechanism [yuan2021agentformer, tang2024hpnet] to better fuse spatiotemporal contextual information.
Recently, alternative generative frameworks have also demonstrated strong performance. Diffusion models [bae2024singulartrajectory, wang2024optimizing], for example, treat trajectory forecasting as an iterative denoising process, generating a diverse set of realistic future paths from an initial random state. Large language models (LLMs) [peng2025lc, yang2025trajectory] are adapted for this task by tokenizing the traffic scene and agent dynamics, leveraging their advanced reasoning capabilities to predict plausible, human-like behaviors with enhanced explainability. For example, by using supervised fine-tuning, LC-LLM [peng2025lc] concurrently predicts the final trajectory and generates natural language explanations for its lane-change intentions.
2.2 Query-Based Trajectory Decoding
@@
Recent end-to-end trajectory prediction models adopt a query-based trajectory decoding paradigm, inspired by the Detection Transformer (DETR) [carion2020end] from the object detection field. Query refers to a set of learnable embedding vectors that serve as placeholders for future trajectory predictions. These queries interact with encoded scene features through attention mechanisms within the Transformer architecture, enabling the model to generate distinct and contextually relevant trajectory forecasts. Various terms have been used to describe this underlying concept, such as ”proposals” in mmTransformer [liu2021multimodal], ”anchor trajectories” in MultiPath++, ”queries” in MTR [shi2022motion], SEPT [lan2023sept], and QCNet [zhou2023query], etc.
@@
Various studies have contributed to improvements in query design. Early studies ultilize predefined queries to inform the model of possible endpoints [gu2021densetnt, zhao2021tnt, shi2022motion]. Recent studies purpose to decode trajctories dynamically with learnable queries [zhou2023query, zhang2024decoupling]. The queries are designed to capture the learned contextual information with cross-attention mechanism and produce multi-modal future trajectories. A recent study [wang2025risk] proposes endpoint-risk-combined intention queries as prediction priors to support risk-aware risk prediction.
2.3 Long-Tail Trajectory Prediction
@@
Long-tail learning addresses the challenge of imbalanced data distributions, where a large portion of the dataset consists of rare or less frequent examples. In the field of trajectory prediction, various strategies have been proposed to tackle this issue, including data augmentation[bahari2022vehicle, li2024pre], loss design [kozerawski2022taming], contrastive learning [makansi2021exposing, wang2023fend], mixture of experts [mercurius2024amend], model ensemble [li2024adaptive], etc.
@@
Input data augmentation techniques are utilized to improve the accuracy and robustness of trajectory prediction, incorporating strategies such as heading rotation, scene flipping, and adding random noise. These strategies have been shown to increase robustness against adversarial patterns in trajectories [zhang2022adversarial, bahari2022vehicle]. The design of synthetic driving data has also demonstrated notable benefits for trajectory prediction [li2024pre]. [ganeshaaraj2025enhancing] altered the input data distribution by an embedding-based clustering technique in a two-phase training scheme.
@@
Contrastive learning (CL) was first proposed to deal with long-tail issuse in trajectory prediction by [makansi2021exposing]. They propose to improve long-tail trajectory prediction performance by forcing similar samples together and pulling dissimilar samples apart in the feature space by a contrastive loss function. [wang2023fend] propose FEND, a feature-enhanced distribution-aware CL framework that ultlizes prototypical contrastive learning (PCL). A following study [zhang2024tract] further integrates contextual scene information in the contrastive learning framework. However, the scene contextual interaction information is not explicitly considered in their contrasitive learning framework. Researchers [yang2024dynamic] also extend the use of contrastive learning by considering subclasses dynamically. [lan2024hi] propose a hierarchical wave-semantic contrastive learning (Hi-SCL) framework, which maintains a collection of feature-enhanced hierarchical prototypes, dynamically steering trajectory samples closer or pushing them farther away.
@@
Contrastive learning typically functions by differentiating between positive and negative sample pairs. In current contrastive learning-based trajectory prediction methods, contrastive sample pairs are often constructed based on difficulty scores [makansi2021exposing] or the clustering of focal agent motion patterns [wang2023fend], frequently overlooking information from interacting agents and the dynamic environment. This approach poses a significant challenge in effectively constructing positive and negative sample pairs that incorporate scene semantics [zhou2024smartpretrain], which in turn hinders the development of a robust contrastive loss function.
In this paper, we avoid the challenge of attempting to apply contrastive learning to consistently consider both interacting- and scene-level clustering. Alternatively, we approach long-tail trajectory prediction through deviation knowledge fusion and dual query-based decoding with loss reweighting. We posit that both individual motion and interaction patterns contribute to the rarity of traffic scenarios and should be carefully designed and incorporated into the model design.
3 Long-Tail Characteristics of Trajectory Data
3.1 Preliminaries
@@
Given a series of $d_{a}$-dimensional observed motion features $\bm{X}\in\mathbb{R}^{N_{a}\times T_{o}\times d_{a}}$ for $N_{a}$ agents, including the target vehicle and its surrounding agents, over a time span $T_{o}$ and high definition (HD) map vectors $\bm{I}\in\mathbb{R}^{N_{m}\times l_{m}\times d_{m}}$ of $N_{m}$ map polylines, our objective is to predict future positions $\bm{Y}\in\mathbb{R}^{T_{f}\times 2}$ of the target vehicle in a certain time horizon $T_{f}$. Specifically, we seek to train a neural network to model the mapping $f(\bm{Y}|\bm{X},\bm{I})$. Additionally, we introduce a long-tail score $S$ to quantify the likelihood of a sample belonging to the tail of the data distribution. On this basis, we seek to enhance the model’s predictive performance in these long-tail scenarios while not affecting the overall performance.
3.2 Long-Tail Definition
@@
Long-tail learning aims to train a deep neural network on a dataset with a long-tailed class distribution, where a small proportion of classes contain a large number of samples, while the majority of classes are represented by only a few samples [zhang2023deep]. In classification tasks, it is relatively straightforward to identify tail samples. However, long-tail trajectory prediction is essentially a long-tail regression task [yang2021delving], where labels (future trajectories) are continuous values. In this case, there are no hard classification boundaries among classes [zhang2023deep]. Previous studies identify tail samples by a difficulty score, which is computed by the (final) displacement error performance of a Kalman filter [makansi2020multimodal, makansi2021exposing] or a trained prediction network [wang2023fend].
@@
In this study, we propose measuring a tail sample with both its difficulty and rarity. We believe that tailed trajectories are not only samples that are challenging to predict, but also tend to exhibit more complex motion patterns and diverse interaction types, making them relatively rare in spatiotemporal distribution. Although these two aspects are not identical, they are closely related and should be considered together in a holistic manner.
To quantify the difficulty score $S_{d}$, we first train a baseline QCNet model. We then use this pre-trained model to perform inference on the entire training set, and the resulting average displacement error for each trajectory is saved as its difficulty score. The pre-calculated scores remain fixed during the subsequent training of our CDKFormer model. The rarity score $S_{r}$ is designed to capture the rarity of the agent motion dynamics, and is composed of spatial rarity $S_{r,s}$ and temporal rarity $S_{r,t}$. Spatial rarity $S_{r,s}$ captures the statistical infrequency of a trajectory’s destination and is determined by the negative log-likelihood of a Gaussian mixture model (GMM) fitted on the 2D final endpoints of the training trajectories. Temporal rarity $S_{r,t}$ captures the rarity of the agent motion dynamic over its entire trajectory. We treat each coordinate as a separate one-dimensional function and apply functional principal component analysis (FPCA) to find their dominant modes of variation. Each trajectory is then represented by a low-dimensional vector of its FPC scores. A second GMM is fitted on the FPC scores, and the temporal rarity is the resulting negative log-likelihood. The rarity score $S_{r}$ is the square root multiplication of spatial rarity $S_{r,s}$ and temporal rarity $S_{r,t}$.
$S_{r}=\sqrt{S_{r,s}\times S_{r,t}}$ (1)
On this basis, samples that are less likely to occur within the learned spatial and temporal distributions are assigned higher rarity scores. Finally, the tail score $S$ is computed as the square root product of difficulty score and rarity score:
@@
The tail score distribution of training samples in Argoverse 2 motion forecasting dataset is demonstrated in Figure 2.
Figure 1: Rarity score distribution. (a) Spatial rarity score distribution. The trajectory endpoints are fitted to a GMM. The score is the negative log-likelihood of an endpoint under this distribution. (b) Temporal rarity score distribution. Calculated from a GMM fitted on the low-dimensional FPCA scores of the full trajectories. (c) Final rarity score distribution. The final score is the square root product of the spatial and temporal rarity scores. All scores are normalized to [0, 1], with higher scores indicating higher rarity. Both GMMs have 10 components, which is selected based on minimizing Bayesian information criterion.
@@
3.3 Long-Tail Characteristics